May Sum has saved $125 000 in an investment account. She will use it to buy an annuity that pays 6.5%/a compounded quarterly. She will receive quarterly payments for the next 25 years. The first payment will be made 3 months from now.

a) What is the quarterly payment she will receive?

b) What is the interest earned over the duration of the annuity?


1
Expert's answer
2021-02-24T06:43:05-0500

a. She will get (0.065/4) of $125000 quarterly

= 0.065/4 × $125000 =$2031.25

b. Total amount A after 25 years = P(1+r/n)nt

= 125000(1+0.065/4)4×25

= 125000(1.01625)100

= $626564.6274

Interest = A-P = $626564.6274 - $ 125000

= $501564.6274